# DeblurDiNAT: A Lightweight and Effective Transformer for Image Deblurring

### v1

@@ 5 Conclusion
In this work, we proposed a novel Transformer, called DeblurDiNAT, and instroduced four key designs to solve existing issues of deblurring Transformers. Firstly, an alternating dilation factor structure is utilized to recognize both local and global blurred patterns. Then, DeblurDiNAT uses a cross-channel learner to modulate the self-attention outputs dynamically, so that cross-channel interactions can be learnt from complex image data. With the aim of efficient feature propagation, a divide and multiply feed-forward network replaces the conventional approach composed of fully connected layers and activation fucntions. In addition, a lightweight gated feature fusion method is created to aggregate multi-scale or same-scale features. Comprehensive experiments demonstrate that our architecture outperforms SOTA models in regard to the efficiency and efficacy balance, and ours exhibits the strongest generalization ability.

### v2

@@ 5 Conclusion
In this work, we proposed a novel Transformer, called DeblurDiNAT, and instroduced four key designs to solve existing issues of deblurring Transformers. Firstly, an alternating dilation factor structure is utilized to recognize both local and global blurred patterns. Then, DeblurDiNAT uses a local cross-channel learner to modulate the self-attention outputs dynamically, so that cross-channel interactions can be learnt from complex image data. With the aim of efficient feature propagation, a divide and multiply feed-forward network replaces the conventional approaches. In addition, a lightweight dual-stage feature fusion method is created to introduce non-linearity and aggregate multi-scale or same-scale features. Comprehensive experiments demonstrate that our architecture outperforms SOTA models in regard to the efficiency and effectiveness balance, and ours exhibits the strongest generalization ability.
Acknowledgments. This work was supported by the Faster GPU cluster built by Texas A&M High Performance Research Computing (HPRC). We also thank NATTEN for the partial support of this work, which is an open-source deep learning library with efficient C++ and CUDA kernels.